From GlueXWiki
Jump to: navigation, search

January 15, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. File:CDCUpdate15Jan2015.pdf (Mike)
  2. FDC update
  3. Electronics
  4. Engineering Chiller Interlock (Dave)
  5. Other


Participants: Naomi (CMU), Eugene, Mike, Dave, Fernando, Chris, Nick, Simon, Cody, Beni, Vlad, and Lubomir (JLab).

CDC update

- Mike presented a report linked above: he and Simon found a parameter in the time-based track reconstruction that was not set correctly. After fixing this the efficiency is much higher (in the >90s% region). Great result! The residuals vs drift time has an offset ~300microns at >100ns which is interpreted as an indication for a different (faster) gas mixture. Note: the results presented are with magnetic field and the tracking was done with time-to-distance as simulated with Garfield for that field. Next step: to extract the time-to-distance from the data. Mike also presented the tool he made for reading/viewing one crate of f125s using pulser.

- CDC efficiency to be discussed on the next meeting.

FDC update

- Vlad used Mike's tool to test the bad FDC channels (see logbook entry). Out of 8 channels, 7 could not see the pulser signal (applied at the input of the pre-amp), i.e. having issues with the pre-amp, cable, connector or the fADC. Chris and Nick will check the 3M connectors for these channels.

- From the alcohol consumption (~0.4l/30days for FDC and ~0.6l/30days for the CDC according to Scot) Lubomir estimated ~0.8% alcohol in the FDC and ~0.4% for the CDC, very crude estimate!!! From the vapor pressure estimations we get the same order of 0.5-1% alcohol. Since the pressure in the alcohol bubbler varies between 7-12psi the alcohol percentage would vary too. As a quick fix we will make the pressure range smaller.

- Lubomir simulated with Garfield the effect of adding 1% propanol to the standard gas of 40/60 Ar/CO2. From the attached plots one can see ~6% decrease in the drift velocity, i.e. opposite to the effect observed in the data. If this is true we would expect much bigger deviation from the standard 40/60 Ar/CO2 mixture than what was estimated from the drift velocity to be 46/54 (see January 8 meeting), i.e. should be about 50/50! We are looking now for a way to measure the gas percentages in the mixture.


- Nick and Chris started modifying the HVBs: adding capacitors, milling retainers. Some dust (like rust) and fibers were found on the caps of the top HVBs - picture linked above.

- Cody: 24 (out 25) new fADCs were tested and only two failed - much better than with the previous production. Cody replaced 10 fADC125 (see logbook entry) following the list of bad channels (with missing bits) given to him by Naomi and Vlad. Will try to fix the bad channels if possible or ship them for repair.


- Dave presented his plan for re-using an old AC power reset module (from Hall B) for shutting down the chiller. It has an AC contactor controlled by the PLC using same logic as for the MPOD interlock. (presentation attached).

- The FDC chiller has been finally shipped for repair.